OpenStudy (anonymous):

what is the product of 5/6 and 2/3? 10/12 5/9 7/9 15/12

OpenStudy (anonymous):

5/6 x 2/3 = (5x2)/(6x3)

OpenStudy (anonymous):

multiply the numerators then multiply the denominators

OpenStudy (anonymous):

find a common denominator by multiplying 6and 3 which will give you 18. then cross multiply 5 and 3 which will give you 15 as well as cross multiply 6 and 2 which will give you 12 and then multiply 15 and 12 divided by 18. (15 x 12 = 180/18= 10). Therefore your answer is 10